3.10 A company purchased new equipment for $250,000 now, that will require annual operations and maintenance (O&M) costs of $30,000 per year for 5 years starting 3 years from now. What is the present worth of these costs if the interest rate is 8% per year?
Answer #1

ANSWER:

Purchase equipment = $250,000

o and m costs = $30,000

i = 8%

Pw = Purchase equipment + o and m costs in year 3(p/f,i,n) + o and m costs in year 4(p/f,i,n) + o and m costs in year 5(p/f,i,n) + o and m costs in year 6(p/f,i,n) + o and m costs in year 7(p/f,i,n)

Pw = 250,000 + 30,000(p/f,8%,3) + 30,000(p/f,8%,4) + 30,000(p/f,8%,5) + 30,000(p/f,8%,6) + 30,000(p/f,8%,7)

Pw = 250,000 + 30,000 * 0.7938 + 30,000 * 0.735 + 30,000 * 0.6806 + 30,000 * 0.6302 + 30,000 * 0.5835

Pw = 250,000 + 23,814 + 22,050 + 20,418 + 18,906 + 17,505

Pw = 250,000 + 102,693

Pw = 352,693

so the present worth is $352,693